OpenStudy (anonymous):

example of quadratic equation that has no (0) real number solutions??

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (solomonzelman):

One question before I do it, are you well familiar with completing the square ?

OpenStudy (bradely):

x^2-4x+53=0 has no real roots

OpenStudy (solomonzelman):

No? Then I can't explain how to do it.

OpenStudy (solomonzelman):

Bradely, in your equation the number can be 5 or more :)

OpenStudy (solomonzelman):

I mean the c>4 in ax²+bx+c=0

OpenStudy (solomonzelman):

I mean when a=1 and b=2 (not necessarily in other cases)

OpenStudy (solomonzelman):

b=4 (not 2)

OpenStudy (tkhunny):

Make b = 0 and you can write them all day. Just make everything else positive. \(4x^{2} + 1 = 0\) \(3x^{2} + 2 = 0\) \(7x^{2} + 9 = 0\)

OpenStudy (solomonzelman):

Yes that's a good way,but I don't like the way the questions with b=0 look :)